OpenStudy (taylovesmatthew):

16 ft= how many inches

jimthompson5910 (jim_thompson5910):

What is 1 foot equal to (in terms of inches)?

OpenStudy (taylovesmatthew):

12

OpenStudy (taylovesmatthew):

i got it thank you!1:)

jimthompson5910 (jim_thompson5910):

1 ft = 12 inches so you'd multiply both sides by 16 16*1 ft = 16*12 inches 16 ft = ??? inches